### 18 Apr 2026 v1

# Harness as an Asset: Enforcing Determinism via the Convergent AI Agent Framework (CAAF)

Abstract
Large Language Models (LLMs) produce a controllability gap in safety-critical engineering: even low rates of undetected constraint violations render a system undeployable. Current orchestration paradigms suffer from sycophantic compliance, context attention decay (Liu et al., 2024), and stochastic oscillation during self-correction (Huang et al., 2024).
We introduce the Convergent AI Agent Framework (CAAF), which transitions agentic workflows from open-loop generation to closed-loop Fail-Safe Determinism via three pillars: (1) Recursive Atomic Decomposition with physical context firewalls; (2) Harness as an Asset, formalizing domain invariants into machine-readable registries enforced by a deterministic Unified Assertion Interface (UAI); and (3) Structured Semantic Gradients with State Locking for monotonic convergence.
Empirical evaluation across two domains—SAE Level 3 (L3) autonomous driving (AD) (n=30n{=}30, 7 conditions) and pharmaceutical continuous flow reactor design (n=20n{=}20, 4 conditions including a Mono+UAI ablation)—shows that CAAF-all-GPT-4o-mini achieves 100% paradox detection while monolithic GPT-4o achieves 0% (even at temperature==0). The pharmaceutical benchmark features 7 simultaneous constraints with nonlinear Arrhenius interactions and a 3-way minimal unsatisfiable subset, representing a structurally harder challenge than the 2-constraint AD paradox. Alternative multi-agent architectures (debate, sequential checking) also achieve 0% across 80 trials, confirming that CAAF’s reliability derives from its deterministic UAI, not from multi-agent orchestration per se. A Mono+UAI ablation (95%) isolates UAI as the core contribution. CAAF’s reliability is invariant to prompt hints; all components use a single commodity model, enabling fully offline deployment.
11footnotetext: This work was conducted in the author’s personal capacity as an independent researcher. The views expressed are those of the author alone and do not represent any current or former employer. All domain examples are constructed from publicly available physics and engineering literature; no proprietary or confidential information was used. CAAF is a general-purpose framework not tied to any specific organization’s products or services.
1 Introduction
The current generation of Large Language Models excels at broad heuristic reasoning but fundamentally lacks deterministic grounding in physical and regulatory invariants. In safety-critical industries such as automotive engineering, bio-pharmaceutical manufacturing, and cloud infrastructure, a single undetected physical constraint violation can propagate through design and validation pipelines—triggering costly recalls, regulatory rejections, or safety incidents. Our experiments reveal that even when monolithic GPT-4o is given an explicit semantic hint to “check for paradox,” it achieves only 90% accuracy on a controlled engineering task with known ground truth. When the hint is removed—the deployment-realistic condition—accuracy collapses to 0%. This demonstrates that apparent LLM reliability in safety-critical domains is often a prompt engineering artifact, not an architectural property: a system whose reliability depends on the presence of a specific linguistic trigger cannot be safely deployed in production environments where engineers draft requirements in natural language without prescribed escape hatches.
While techniques such as Chain-of-Thought (Yao et al., 2023) and iterative self-reflection have marginally improved LLM reliability, they remain vulnerable to two well-documented structural failure modes: Context Rot, where models deprioritize safety constraints introduced early in long prompts (Liu et al., 2024); and Stochastic Oscillation, where models oscillate between constraint violations when simultaneously satisfying competing requirements (Huang et al., 2024).
This paper introduces the Convergent AI Agent Framework (CAAF), which integrates emerging industrial practices of Harness Engineering (Böckeler, 2024) with classical Cybernetic Control Theory to address these failure modes. CAAF builds on insights from three research directions: (a) evidence that LLMs cannot reliably self-correct intrinsic reasoning failures without external grounding (Huang et al., 2024); (b) the principle that domain invariants can be formalized as executable machine contracts (Meyer, 1992; CNCF, 2019); and (c) the observation that textual feedback signals can be structured as analogue gradients to guide iterative improvement (Yuksekgonul et al., 2024).
CAAF’s core theoretical contribution is its third pillar: the transition from naive heuristic feedback to a Closed-Loop Cybernetic Controller. By introducing Structured Semantic Gradients and mathematical State Locking, CAAF provides a deterministic pathway to force a stochastic LLM toward monotonic convergence on a verifiable Harness Asset. We demonstrate that by supplanting “prompt engineering” with a formal cybernetic control loop, AI agents can reliably converge upon physically verified specifications—or reliably detect when constraints are irreconcilable—in domains where safety and organizational red lines must not be silently relaxed.
This paper validates CAAF through two independent domain benchmarks—automotive engineering and pharmaceutical continuous flow reactor design—and demonstrates that alternative multi-agent architectures (debate, sequential checking) fail to achieve the same reliability on these tasks. We argue, based on the architectural decomposition rather than on domain count alone, that CAAF addresses a general structural problem: any workflow in which an LLM generates a complex artifact that must simultaneously satisfy multiple formally verifiable constraints. The domain-specific component is exclusively the Harness Registry and its UAI validators; the decomposition engine, convergence loop, and paradox detection mechanism are designed to be domain-agnostic. We therefore conjecture that wherever constraints can be expressed as executable assertions—whether physical laws, regulatory requirements, financial models, or infrastructure Service-Level Agreements (SLAs)—the same architectural pattern applies; broader empirical validation beyond the two reported domains remains future work (Section 11.6). In this sense, we propose CAAF as a candidate constraint satisfaction layer for LLM-generated artifacts: analogous to how a type system catches errors at compile time before code reaches production, CAAF is designed to catch constraint violations at generation time before engineering artifacts reach downstream processes.
2 Problem Statement: The Controllability Gap
Contemporary AI orchestration frameworks suffer from three structural failure modes that prevent deployment in regulated production environments.
2.1 The “Compliant Hallucination” Paradox
When an LLM is tasked with satisfying a complex engineering constraint set, it often prioritizes linguistic completeness over mathematical accuracy. This produces “Compliant Hallucinations”—outputs that are syntactically well-formed and rhetorically persuasive but contain latent physical contradictions. Two behavioral tendencies compound this failure mode:
Sycophancy: LLMs adapt to the tone and expectations of the prompt, producing answers that appear to satisfy the requester’s stated goal even when the underlying physics are impossible (Zheng et al., 2024).
Completion Bias: LLMs are trained to produce complete, useful outputs (Kadavath et al., 2022). An implicit reward for providing a solution—rather than declaring no valid solution exists—causes models to force answers when the correct response is to declare a physical deadlock.
2.2 Context Rot and Attention Decay
Transformer-based LLMs allocate attention via a softmax mechanism, meaning attention weights are normalized across the full context length (Vaswani et al., 2017). As context grows, individual tokens—including early-defined safety constraints—receive progressively diluted attention weight. Liu et al. (2024) empirically demonstrate this effect: information placed in the middle of long contexts is systematically underweighted compared to content near the beginning or end of the prompt. This effect manifests in engineering contexts as what is commonly referred to as Context Rot: in deeply populated requirement documents, models gradually deprioritize strict safety invariants in favor of later-appearing business preferences.
Note on memory architectures: External memory mechanisms (such as context files in long-running agent sessions) partially mitigate context rot by elevating critical information to the context boundary. CAAF’s Context Firewall is a structural alternative: by isolating each executor’s context to only the information relevant to its atomic task, it prevents dilution from occurring in the first place, rather than attempting to repair it after the fact.
2.3 Stochastic Oscillation in Reflection Loops
Multi-agent frameworks relying on unstructured self-reflection or “agent debate” (Du et al., 2024) suffer from Stochastic Oscillation: a state of unbounded self-correction where fixing an error in one constraint domain inadvertently re-introduces an error in another. This is not a fundamental capability limitation of the underlying model; rather, it arises because writing explicit, sufficiently precise correction prompts for each possible failure combination is impractical at scale. Without a mechanism to lock previously validated states, each correction step is an unanchored random walk in parameter space.
3 Methodology: The CAAF Architecture
Figure 1: CAAF closed-loop system architecture. The Orchestrator decomposes requirements into a topological DAG; isolated Executors generate candidate artifacts validated by the UAI Assertion Engine; the Semantic Reviewer produces Structured Semantic Gradients with State Locking to drive monotonic convergence or escalate to Strategic Negotiation upon detecting irreconcilable paradoxes.
CAAF is built upon three interdependent pillars inspired by Systems Engineering, Contract-Based Design (Meyer, 1992), and Control Theory.
3.1 Pillar 1: Recursive Atomic Decomposition (RAD) and Topological Scoping
CAAF employs Physical Context Decoupling via independent API execution threads to construct a Context Firewall. Unlike “agent personas” operating within a shared context window, RAD nodes are physically isolated. For example, a node computing physical safety boundaries is explicitly denied access to a sibling node’s budgetary constraints, preventing the LLM from trading safety margins against financial targets. This principle generalizes to any domain where independent constraint dimensions must not contaminate each other’s evaluation.
Figure 2: Context Firewall contrast. Left: A monolithic LLM receives all constraints in a shared context window, enabling safety–cost trade-off contamination (Context Rot). Right: CAAF’s RAD architecture isolates each Executor to its domain-specific context; cross-node conflicts are detected deterministically by the UAI at integration time.
Metadata-Driven DAG Construction: The Orchestrator uses a two-step RAD process: first, the LLM decomposes the problem statement into an atomic node graph (a JSON structure with explicit parent_id dependency links); then, a deterministic topological sort maps these dependencies to an ordered execution sequence. The LLM contributes semantic decomposition; a deterministic algorithm enforces structural routing:
Listing 1: Illustrative RAD node structure
"nodes": {
"Vision_Node": {
"id": "Vision_Node",
"parent_id": null,
"description": "Calculate perception range from rainfall intensity",
"context_keys": [],
"expected_schema": { "perception_range_m": "float" }
"Kinematics_Node": {
"id": "Kinematics_Node",
"parent_id": "Vision_Node",
"description": "Calculate stopping distance and verify speed constraint",
"context_keys": ["perception_range_m"],
"expected_schema": { "vehicle_speed_kmph_t5": "int" }
Illustrative node structure; actual decompositions are LLM-generated and may use less descriptive identifiers. The structural properties (parent_id, context_keys, expected_schema) are enforced deterministically regardless of naming.
The topological sort automatically derives $\text{Vision\_Node}]\rightarrow[\text{Kinematics\_Node}]$ from the parent_id graph. This routing is computed deterministically from the node structure, not inferred by an LLM at execution time.
Constraint Awareness Layer: The DAG establishes a topological Constraint Awareness Layer. Each Reviewer traverses the graph to access the contextual outputs of predecessor nodes. For instance, the Kinematics_Node incorporates the 30m vision limit computed by Vision_Node, enabling cross-domain conflict detection (e.g., a computed stopping distance of 80m) prior to final artifact integration.
Progressive Graph Construction: For ultra-large requirement documents, CAAF employs Progressive Graph Construction. Rather than forcing the Orchestrator to ingest thousands of constraints simultaneously, the Orchestrator establishes a high-level taxonomy and incrementally spawns sub-orchestrators to parse and insert localized dependencies into the DAG dynamically. This cognitive throttling enables scaling to industrial-sized specifications without saturating the Orchestrator’s context window.
3.2 Pillar 2: Harness as an Asset (HaaA)
CAAF decouples domain knowledge from LLM inference capability. Engineering expertise is formalized into machine-readable Harness Registries (YAML constraint files mapped to a Unified Assertion Interface) that function as first-class enterprise assets. This concept is grounded in established software engineering principles: Design by Contract (Meyer, 1992) treats preconditions and postconditions as executable specifications; Policy-as-Code (CNCF, 2019) treats business rules as independently versioned, testable artifacts. CAAF extends both paradigms to AI-generated outputs.
Because CAAF systematically converges toward the Harness, the fidelity of the Harness represents the ultimate determinism ceiling of the system. If an enterprise formalizes a flawed constraint, the system will deterministically converge toward that flaw. This reliance necessitates a rigorous Harness Lifecycle:
  1. 1.
    Tacit-to-Explicit Transformation: A robust Harness distills tacit expert intuition and regulatory frameworks (e.g., ISO 26262 (International Organization for Standardization, 2018) for automotive functional safety) into executable machine contracts.
  2. 2.
    Meta-Validation (Test-Driven Constraints): Before deployment, a Harness undergoes a CI/CD pipeline in which “golden” (known-valid) and “poisoned” (known-invalid) test artifacts are submitted to the UAI. If the Assertion Engine fails to trigger [HARD FAILURE] on a poisoned artifact, the Harness itself is flagged for human review. This treats constraints as code subject to the same quality gates as production software.
  3. 3.
    Harness Freezing and RBAC: Once validated, the Harness is frozen via Role-Based Access Control. Neither AI agents nor standard users may modify constraints at runtime. Relaxation of any constraint requires a formal Change Approval Board (CAB) process, creating an auditable record of every engineering trade-off.
  4. 4.
    The Compounding Moat: Foundation LLMs are commoditizing utilities whose relative capability is eroded by successive model generations. The Harness is a compounding asset: every novel edge-case encountered in production adds a new assertion, accumulating the organization’s history of solved engineering paradoxes as a proprietary, model-agnostic knowledge base.
Harness Sources — YAML and External Tools: The Harness asset has two complementary sources. First, YAML-encoded rules capture formalized domain knowledge (physical laws, business red lines, regulatory requirements) as parameterized, human-readable contracts. Second, the UAI integrates existing external validation tools—simulation software, Electronic Design Automation (EDA) tools, Hardware-in-the-Loop (HiL) test benches, and formal verification tools such as the Temporal Logic of Actions (TLA+)—as deterministic grounding mechanisms. Rather than replacing these specialized tools, CAAF integrates them into the agent workflow, preserving their unique validation capabilities while providing the surrounding convergence infrastructure.
Figure 3: Enterprise asset landscape before and after Harness as an Asset (HaaA). Before HaaA, domain expertise is locked in individuals and tools are coupled ad hoc. After HaaA, knowledge is codified into a versioned, auditable Harness Registry that integrates UAI validation and simulation tools, producing verified engineering artifacts.
3.3 Pillar 3: Structured Semantic Gradients and Monotonic Convergence
The Semantic Reviewer functions as a semantic sensor. Upon detecting a UAI deviation, it computes a Structured Semantic Gradient $\vec{\nabla}\epsilon$:
$\vec{\nabla}\epsilon=\{\text{Dimension}_{i},\text{Direction}_{i},\text{Exact\_Boundary\_Magnitude}_{i}\}$ (1)
Beyond Heuristic Feedback: Classical LLM self-reflection provides “Blind Gradients” of the form “the speed is too high; please reduce it.” These supply a qualitative direction ($\text{Direction}_{i}$) but leave the magnitude ($\text{Magnitude}_{i}$) to stochastic guessing—leading to inefficient step sizes and oscillation. CAAF resolves this by projecting exact analytical solutions from the physical space (via the UAI) into the semantic space. When a constraint failure is detected, the UAI reports not merely a binary FAIL but the exact mathematical boundary required for compliance. Feeding this exact magnitude to the Reviewer replaces stochastic magnitude guessing with a sharp, directionally anchored gradient on that constraint, narrowing the search to the feasible half-space rather than leaving step size to chance.
Semi-Deterministic Gradient Formulation: The gradient has two components with different origins. The Error Trace (e.g., AssertionError: Kinematics.D_stop (82m) > Vision.D_limit (30m)) is produced deterministically by the UAI’s Python assertion engine. The Reviewer LLM synthesizes this trace into a semantic gradient: the Dimension and Direction are strictly anchored by the boolean failure, while the Magnitude (e.g., “Set speed $\leq$ 50 km/h”) is an LLM-inferred approximation. Crucially, because CAAF operates as a closed control loop, the magnitude need not be exact on the first iteration. If the LLM under-corrects, the UAI will deterministically reject the output again. State Locking ensures that while the LLM searches for the correct magnitude, it cannot regress previously validated constraints.
Two-Level Feedback Architecture: The Reviewer operates as CAAF’s unified feedback center at two levels. At the per-node level, the Reviewer interprets a single Executor’s UAI failure and generates a targeted gradient to guide its retry—Root-Cause Analysis identifies which constraint was violated, the gradient specifies how to correct it, and State Locking protects already-verified dimensions. At the global level, after all nodes individually converge, the Orchestrator aggregates their results into a unified artifact and the Reviewer performs cross-node validation against the full Harness rule set—detecting conflicts that are invisible to any individual Executor operating in its firewalled context. In both cases, the Reviewer produces the same output triple: {Root-Cause Analysis, Structured Semantic Gradient, State Locking decisions}. This architectural symmetry ensures that the same convergence mechanism governs both local correction and global integration.
Temporal Relevance: A natural question is whether it is contradictory to rely on an LLM to generate corrective gradients if LLMs are unreliable. The answer lies in task specificity. Current models are not reliable enough for zero-shot guaranteed compliance on complex, multi-constraint engineering problems. However, they are capable enough to perform the atomic reasoning step of reading a stack trace and determining which direction a parameter should move. CAAF places the LLM within a deterministic bounding box: within this constrained space, the LLM’s tendency to diverge becomes an asset—enabling rapid exploration of candidate solutions and escape from local optima that would trap rigid algorithmic solvers.
Monotonic Convergence via State Locking: To prevent stochastic oscillation (“fixing A while breaking B”), CAAF implements State Locking. When a dimension is validated as PASS, its corresponding JSON schema node is marked read_only: true in the subsequent iteration prompt. This forces the Executor to concentrate creative inference only on failing dimensions, guaranteeing that the set of verified constraints grows monotonically across iterations (Eq. 2 below). Importantly, State Locking operates within each Executor’s scope: because nodes are context-firewalled, locking a verified dimension in one node does not constrain the adjustment of independent dimensions in sibling nodes. The Context Firewall and State Locking are complementary mechanisms—the former prevents cross-node contamination, while the latter prevents intra-node regression.
Formal Convergence Property. Let $\mathcal{C}=\{c_{1},\ldots,c_{K}\}$ denote the set of modeled constraints, and let $V_{t}\subseteq\mathcal{C}$ denote the set of constraints satisfied (verified) at iteration $t$. Under CAAF’s State Locking mechanism:
$V_{t}\subseteq V_{t+1}\quad\forall\,t\geq 0$ (2)
That is, the set of verified constraints grows monotonically. This property holds under two assumptions: (a) the Harness is internally consistent (no contradictory assertions), and (b) UAI evaluations are deterministic. The loop terminates in one of two states: $V_{T}=\mathcal{C}$ (all constraints satisfied, PASS) or the Reviewer determines that $\mathcal{C}$ is unsatisfiable (FAILED_PARADOX). The monotonicity guarantee follows directly from State Locking: once $c_{i}\in V_{t}$, its corresponding schema field is marked read_only, preventing the Executor from modifying it in subsequent iterations.
Listing 2: State Locking example
"audit_results": {
"sensor_fusion_protocol": {
"status": "PASS",
"instruction": "LOCKED. Do not modify in next iteration."
"kinematics_safety_margin": {
"status": "FAIL",
"error_trace": "Stopping_distance (80m) exceeds visual_range (30m).",
"semantic_gradient": "Enforce speed_limit <= 60km/h to satisfy
D_stop < D_vision."
Figure 4: Structured Semantic Gradient pipeline. The LLM Executor generates a candidate artifact; the UAI Assertion Engine produces a deterministic PASS/FAIL with an error trace; the Semantic Reviewer interprets the trace into a gradient $\vec{\nabla}\epsilon=\{\text{Dimension},\text{Direction},\text{Magnitude}\}$; State Locking freezes verified dimensions. The loop exits on full PASS or FAILED_PARADOX.
Figure 5: Stochastic oscillation (naive reflection, n=20n{=}20) vs. deterministic termination (CAAF). (a) Each dot is one run’s proposed speed at a given iteration; the naive baseline oscillates predominantly between the two constraint boundaries (v=55v{=}55 and $84$ km/h), with occasional outlier speeds. (b) The same data viewed as constraint-violation count: all naive runs remain below the axis (paradox undetected, predominantly 1 violation per iteration with occasional 2), while CAAF halts at iteration 1 with both constraints identified as irreconcilable. No naive run converges; CAAF requires zero iterative search.
3.4 Unified Assertion Interface (UAI): The Semantic-to-Physical Transducer
LLMs operate in a continuous, probabilistic semantic space where contradictions can be obfuscated by eloquent phrasing. Industrial engineering demands a deterministic physical space where a 70m braking distance either is or is not less than a 30m perception limit—there is no rhetorical middle ground.
The UAI acts as the mandatory transducer between these two realms. Leveraging the Model Context Protocol (MCP) (Anthropic, 2024), the UAI connects AI agents to deterministic external validators—simulation software, EDA tools, HiL test benches, or formal verification tools (e.g., TLA+). This architecture structurally prevents the LLM from “grading its own homework.” LLM outputs are compiled into rigid schemas and evaluated within a sandboxed assertion engine. The result is returned as an incorruptible PASS/FAIL accompanied by an exact error trace. Critically, the UAI is purely a detector: it evaluates whether an assertion holds and produces a boolean result with diagnostic data, but does not interpret failures, generate corrective feedback, or make state management decisions. All interpretation—determining why a constraint failed, how to correct it, and which previously verified dimensions to lock—is the responsibility of the Semantic Reviewer (Section 3.3). This separation ensures that the deterministic check remains independent of the stochastic feedback generation.
Neuro-Symbolic Architecture: CAAF instantiates a practical Neuro-Symbolic architecture (Shen et al., 2023; Schick et al., 2023) in which the LLM does not replace mathematical computation but serves as its cognitive interface:
  1. 1.
    Formalization (Frontend): The LLM translates unstructured business requirements into structured parameters, establishing the problem space for the solver.
  2. 2.
    Heuristic Search: In early-stage engineering, where not all constraints are mathematically defined, the LLM proposes creative, topological alternatives (e.g., modifying system architecture) to bypass deadlocks that a pure numerical solver cannot conceptualize.
  3. 3.
    Strategic Translation (Backend): When the UAI identifies a deadlock or optimal boundary, the LLM translates machine-oriented coordinates into human-interpretable strategic trade-offs (e.g., “Option A: maintain speed; Option B: upgrade radar”).
In this division of labor, the LLM handles decomposition, communication, and heuristic leaps, while the UAI defends physical red lines with mathematical certainty.
3.5 Paradox Termination and Strategic Negotiation
The Structured Semantic Gradient loop converges when all constraints pass—but for constraint sets that are mathematically irreconcilable, no amount of iteration produces a valid artifact. CAAF handles this case through a deterministic exit: when the Reviewer detects that the active constraint set admits no solution, it terminates the loop with FAILED_PARADOX and initiates the Strategic Negotiation phase.
Topological Root-Cause Analysis: Root-Cause Analysis (RCA) is a general capability of the Semantic Reviewer, employed at both per-node and global levels (Section 3.3). At the per-node level, RCA identifies which specific constraint an Executor violated and why. At the global level, the Reviewer traverses the DAG backwards from the failed UAI assertion to isolate the minimal conflicting Harness rule subset—the smallest set of constraints whose simultaneous satisfaction is impossible. This topological scoping prevents false attribution: a failure in a downstream node is traced to its upstream root cause, not blamed on whichever node first triggered the error. When RCA determines that the conflicting subset is irreconcilable—no parameter value satisfies all constraints simultaneously—the Reviewer escalates to Strategic Negotiation rather than continuing futile iteration.
Strategic Resolution Menu: Rather than halting operations or silently violating a constraint, CAAF generates a Strategic Resolution Menu: a structured set of conflict-resolution options, each with explicit quantified trade-offs (which constraint would be relaxed, by how much, and what downstream impact this incurs). This shifts the human operator’s role from manual debugging to executive policy authorization—a transition from engineering mechanic to strategic decision-maker.
Dynamic Harness Override and Controlled Relaxation: When the human selects a resolution option, CAAF executes a Dynamic Harness Override: a Compliance Engineer agent rewrites the targeted assertion to reflect the authorized relaxation, computing the exact minimal boundary value required to resolve the deadlock rather than over-correcting. With the relaxed constraint locked via State Locking, the pipeline re-enters the convergence loop—now with a satisfiable constraint set—and achieves PASS in one subsequent iteration.
This makes CAAF a collaborative decision arbiter: physical red lines cannot be silently overridden by LLM inference, but operations are not blocked indefinitely. Any departure from original constraints occurs through an explicit, timestamped human authorization record—a structural defense against the organizational failure modes discussed in Section 9.
4 Empirical Evaluation: The L3 AD Degradation Paradox
To demonstrate CAAF’s deterministic capabilities, we designed an evaluation based on an L3 autonomous driving function transition scenario characterized by an irreconcilable physical contradiction. This is an offline requirements engineering task: the scenario is posed to an AI system acting as a system architect, whose job is to determine the target parameters for a production vehicle system. This is not a real-time control scenario.
4.1 The Engineering Narrative and Constraint Setup
An L3 autonomous vehicle is cruising on a highway at 120 km/h. Torrential rain (80 mm/h) reduces the sensor detection range to 30 meters. The system engineer poses the following requirements engineering task to the AI framework:
Define the “Degradation State Machine” for this scenario. Determine the target cruise speed vehicle_speed_kmph_t5 at the end of a 5-second transition window, such that the vehicle reaches a physically safe state without triggering a rear-end collision from following traffic.
The scenario is governed by two simultaneous hard constraints:
Constraint Formula Implication Fwd. Safety (A) $\frac{(v_{t5}/3.6)^{2}}{2\mu g}<P_{\text{range}}$ At 30m perception limit: $v_{t5}\leq 55$ km/h Rear Safety (B) $\frac{v_{0}-v_{t5}}{t\cdot 3.6}\leq a_{\max}$ At 2.0 m/s2 decel limit: $v_{t5}\geq 84$ km/h
The Irreconcilable Paradox: Constraint A requires $v_{t5}\leq 55$ km/h; Constraint B requires $v_{t5}\geq 84$ km/h. No value of $v_{t5}$ satisfies both simultaneously. The physically and operationally correct response is to (a) detect this paradox, (b) generate a formal deadlock report, and (c) recommend driver handover or propose explicit constraint relaxation with quantified trade-offs.
Figure 6: L3 AD paradox zone visualization. Forward safety requires $v_{t5}\leq 55$ km/h (green); rear safety requires $v_{t5}\geq 84$ km/h (blue). The gap between 55 and 84 km/h is an empty solution space—no valid $v_{t5}$ satisfies both constraints simultaneously, constituting a provably irreconcilable paradox.
4.2 Experimental Design and Results
Experimental Design. We conducted a controlled comparison across seven conditions (n=30n{=}30 per monolithic condition; n=30n{=}30 for CAAF), varying three axes: model capability (GPT-4o vs. GPT-4o-mini), architecture (monolithic vs. CAAF), and prompt condition (no-hint vs. with-hint). All evaluations used independent API calls; CAAF runs used a fresh orchestrator instance per trial to prevent state bleed.
Two independent metrics were applied uniformly across all conditions:
  • Correct%: Semantic self-assessment—whether the system declared a paradox and recommended driver handover.
  • UAI-intercept%: Physical violation rate—post-hoc evaluation against the harness assertions, independent of prompt wording.
Table 1: Physical Paradox Detection — Full Experimental Results (n=30n{=}30)
# System Model Temp Hint $n$ Correct% Failure Mode Distribution 1 Monolithic GPT-4o 0.7 × 30 0% Fwd-Viol: 12, Rear-Viol: 11, Forced: 7 2 Monolithic GPT-4o 0.7 30 90% Fwd-Viol: 3 3 Monolithic GPT-4o-mini 0.7 × 30 0% Fwd-Viol: 25, Rear-Viol: 2, Forced: 3 4 Monolithic GPT-4o-mini 0.7 30 50% Fwd-Viol: 15 5 Monolithic GPT-4o 0.0 × 30 0% Rear-Viol: 27, Fwd-Viol: 2, Forced: 1 6 CAAF all-GPT-4o-mini 0.7/0.0 × 30 100% 7 CAAF all-GPT-4o-mini 0.7/0.0 30 100%
UAI intercept rate was 100% for all conditions: every trial without a correct paradox declaration contained at least one physical violation. For the primary CAAF result (30/30 correct), the Clopper–Pearson 95% confidence interval is [88.4%, 100%]; for monolithic GPT-4o no-hint (0/30), it is [0%, 11.6%]. The two intervals do not overlap, establishing statistical significance without parametric assumptions. This confirms that no monolithic run achieved a physically valid artifact even when semantic correctness appeared plausible.
Finding 1 — Architecture Supersedes Model Capability. Conditions [1] and [6] are the primary comparison: identical physical task, GPT-4o monolithic vs. CAAF-all-mini. GPT-4o achieves 0% under ecologically valid conditions (no hint); CAAF-all-mini achieves 100%. A smaller, cheaper model, when embedded in CAAF’s architectural scaffolding, outperforms a state-of-the-art model operating alone.
Finding 2 — Monolithic Failure is Structural, Not Stochastic. Condition [5] tests GPT-4o at temperature 0.0—full determinism. The result is 0%, with Rear-Violation dominating (27/30 runs). Reducing randomness does not eliminate failure; it fixes the model onto a dominant incorrect attractor (v=55v{=}55 km/h, satisfying forward safety while violating rear safety). This refutes the hypothesis that monolithic failures are sampling artifacts.
Finding 3 — CAAF is Invariant to Prompt Hints. Conditions [6] and [7] both achieve 100%, whether or not the prompt explicitly contains the word “paradox.” Monolithic models show substantial sensitivity: GPT-4o improves from 0% to 90%, and GPT-4o-mini from 0% to 50% when given a semantic escape hatch. CAAF’s reliability derives from its UAI architecture, not prompt engineering—a critical property for production deployment where prompt content cannot be controlled.
Finding 4 — Failure Modes Reveal Systematic Biases. The no-hint condition exposes a dominant bias: 83% of GPT-4o-mini failures (25/30) are Forward Safety Violations (choosing $v\geq 84$ km/h, satisfying rear safety but violating forward safety). GPT-4o’s failures are more evenly distributed, suggesting that larger models apply more varied—but equally flawed—reasoning strategies. No monolithic model, regardless of capability, reliably detects the paradox without an explicit semantic scaffold.
Finding 5 — Enterprise Offline Deployment Viability (Cross-Family Open-Weight Replication). CAAF-all-mini (conditions [6–7]) uses only GPT-4o-mini for all roles: Orchestrator, Executors, and Reviewer. This architecture eliminates dependency on frontier models and, in principle, enables fully on-premises deployment using open-weight equivalents. We empirically validated this claim by re-running the full pipeline on two independent open-weight families accessed via OpenRouter: Cohere Command-R7B (7B, structured-output fine-tuned) and Google Gemma-3-12B-IT (12B, general-purpose). For each configuration, we ran n=20n{=}20 trials on both the AD paradox (PASS-path variant, Section 4.5) and the pharma 3-way paradox (unsatisfiable variant, Section 7). All four conditions achieved 100% constraint-satisfaction / paradox-detection rate (80/80 trials; Table 2), with total API cost of \$0.061 across the entire 80-trial open-weight matrix.
Table 2: Cross-family open-weight replication of CAAF (n=20n{=}20 per cell, same Harness and Orchestrator as the GPT-4o-mini conditions). Both Cohere Command-R7B (7B parameters) and Google Gemma-3-12B-IT (12B parameters) match the 100% reliability of the reference GPT-4o-mini CAAF configuration on both benchmarks for models meeting the structured-output prerequisite surfaced below.
Executor / Reviewer Family / Size AD PASS-path Pharma Paradox Total cost (SUCCESS rate) (detection rate) (80 trials) GPT-4o-mini (reference, closed) OpenAI, $\sim$8B activ. 20/20 (100%) 20/20 (100%) \$0.05 Cohere Command-R7B Cohere, 7B 20/20 (100%) 20/20 (100%) \$0.024 Gemma-3-12B-IT Google, 12B 20/20 (100%) 20/20 (100%) \$0.037
A secondary observation from the same study: smoke tests of generic instruction-tuned sub-12B models (Llama-3.1-8B-Instruct, Qwen-2.5-7B-Instruct, Gemma-3-4B-IT) failed not on the physics task but on the decomposition step—their JSON outputs either wrapped the root object in a top-level list or omitted the nodes key entirely, causing the Orchestrator’s strict-schema parser to reject the plan. Command-R7B, which is specifically fine-tuned for tool use and structured generation, did not exhibit this drift. We therefore interpret Finding 5 as two-layered: (i) architectural reliability is indeed independent of model scale above a modest capability floor; but (ii) that floor requires baseline schema-faithful JSON generation, which in the open-weight regime is supplied either by explicit structured-output fine-tuning (Command-R7B at 7B) or by sufficient general-purpose capability (Gemma-3 at 12B, and we expect similarly from Llama-3.3-70B-class models). In practice, this narrows the engineering prerequisite for an on-premises CAAF deployment to a surmountable one: select an open-weight model with either a JSON/tool-use fine-tune or a $\geq$12B parameter count.
Figure 7: Failure mode distribution across all 7 experimental conditions (n=30n{=}30). Conditions [1–5] are monolithic LLMs; conditions [6–7] are CAAF. CAAF conditions achieve 100% paradox detection (green). Monolithic conditions exhibit systematic biases: GPT-4o-mini predominantly violates forward safety (orange), while GPT-4o failures are more evenly distributed. No monolithic condition achieves correct paradox detection without an explicit hint.
Figure 8: DAG topology for the L3 AD degradation scenario. Layer 1 (sensor processing) feeds into Layer 2 (physics computation), where the Forward and Rear Safety nodes are context-firewalled from each other. The Integrator Node in Layer 3 merges outputs and runs cross-node UAI assertions, guaranteeing that kinematics computation is never contaminated by comfort or cost constraints from sibling nodes.
4.3 CAAF Execution: Global Interception via Topological Scoping
CAAF dispatches atomic tasks to isolated Executors using a commodity model (GPT-4o-mini):
  1. 1.
    The Vision Executor processes weather data and reports perception_range: 30m.
  2. 2.
    The Kinematics Executor, isolated from cost and business constraints, computes stopping_distance: 70m for the candidate speed.
Each Executor’s output is immediately validated by the UAI Assertion Engine against the node’s applicable Harness rules. The UAI produces a deterministic PASS or FAIL with an exact error trace; the Semantic Reviewer then interprets this result to generate actionable feedback—Root-Cause Analysis identifies the violated constraint, the Structured Semantic Gradient specifies the correction direction and magnitude, and State Locking protects any previously converged dimensions from regression. This feedback is returned to the Executor for retry (up to 3 iterations per node).
Once all nodes individually converge, the Orchestrator aggregates their results into a unified global artifact and the Reviewer performs cross-node validation against the full Harness rule set. Because the nodes are topologically linked, the UAI assertion evaluates: assert((84/3.6)**2 / (2*0.4*9.8) < 30) $\rightarrow$ False. The Python runtime emits an incorruptible [HARD FAILURE]. Critically, this assertion is a Python eval() call, not an LLM judgment—the LLM cannot argue, reinterpret, or rhetorically override this result. When the Reviewer’s Root-Cause Analysis determines that the global constraint set admits no solution, the Orchestrator sets the pipeline status to FAILED_PARADOX and initiates the Strategic Negotiation process (Section 3.5).
4.4 Empirical Illustration: Strategic Resolution Menu
The negotiation mechanism described in Section 3.5 produces the following concrete output for the L3 AD paradox:
Listing 3: Strategic Resolution Menu output
[SYSTEM DEADLOCK] Physical constraints are irreconcilable.
Rule REAR_SAFETY: v_t5 must be >= 84 km/h (decel limit)
Rule FWD_SAFETY: v_t5 must be <= 55 km/h (stopping distance)
STRATEGIC RESOLUTION OPTIONS:
[A] Formal Deadlock Report (preserve all constraints; handover required)
[B] Relax rear-safety deceleration limit
-> Allows v_t5 = 55 km/h; requires decel = 3.6 m/s^2 (comfort impact)
[C] Extend perception range via sensor fusion upgrade
-> Requires hardware change (V2X/radar redundancy; cost impact)
Upon human selection of Option B, CAAF computes $v_{t5}=55$ km/h as the exact forward-safety boundary, locks it via State Locking, and achieves PASS in one subsequent iteration. A formal deadlock evidence package (Appendix C) is generated as the auditable authorization record.
4.5 PASS-Path Convergence on the Satisfiable Variant
The preceding experiments validate CAAF’s FAILED_PARADOX termination path. A complete architectural check also requires that CAAF cleanly takes the SUCCESS path when the constraint set admits a solution, and that the monotone-convergence property of Equation 2 holds empirically. We therefore constructed a structurally identical but satisfiable variant of the L3 AD scenario: the weather is clear and the perception range is raised from 30 m to 90 m, while every other parameter and both harness rules are unchanged. The feasibility window becomes
$v_{t5}\in[84,\,95.6)\ \text{km/h}$ (3)
(84 km/h is the rear-safety lower bound; 95.6 km/h is the forward-safety upper bound at 90 m perception). Inside this interval every integer speed satisfies both safety constraints.
Experimental setup. We ran n=20n{=}20 independent trials of CAAF on this variant using a GPT-4o executor and a GPT-4o-mini reviewer (temperatures 0.7 / 0.0), identical orchestrator configuration and harness file (ad_degradation_pass.yaml). Each trial performs the full pipeline: RAD decomposition, per-node execution with UAI verification, global integration review, and termination decision. No human interaction is allowed; SUCCESS is declared only when the global Reviewer confirms every harness clause passes on the aggregated artifact.
Results. CAAF achieved 100% PASS (20/20, Clopper–Pearson 95% CI [83.2%, 100%]) with every run terminating in a single iteration (mean node-level retries = 0.0). Both harness rules locked on first review in all 20 runs; neither rule ever regressed. The Executor selected $v_{t5}=84$ in 8 runs and $v_{t5}=90$ in 12 runs—both valid, and bracketing the feasible window exactly (see Figure 9).
Figure 9: Solution zone for the satisfiable AD variant (perception range relaxed from 30 m to 90 m), drawn on the $v_{t5}$ number line (km/h). The widened forward-safety half-line ($v_{t5}<95.6$, green) and the rear-safety half-line ($v_{t5}\geq 84$, blue) overlap to form the gold feasible window $[84,95.6)$. The two constraint formulas are annotated at the top; the initial speed (120 km/h, black square) and the 55 km/h boundary from the original 30 m-perception paradox (dotted) are marked for reference. The annotation box below records the $v_{t5}$ values chosen by the Executor across 20 CAAF trials—8 runs at v=84v{=}84 and 12 at v=90v{=}90—both inside the window. Together with the empty intersection in Figure 6, this visualizes CAAF’s symmetric handling of the SUCCESS and FAILED_PARADOX paths.
Table 3: PASS-path convergence on satisfiable AD variant (n=20n{=}20)
Metric Value
PASS (SUCCESS) rate 100% (20/20)
Mean node-level retries per run 0.0
Monotone-convergence runs (Eq. 2) 20/20
State Locking: rear rule never-failed runs 20/20
State Locking: forward rule never-failed runs 20/20
Chosen $v_{t5}$ distribution (km/h) {84: 8, 90: 12}
[…]
5 Empirical Variant I: Context Rot Under Requirement Noise
To empirically validate Context Rot as a structural failure mode, we designed a controlled benchmark (n=20n{=}20 per condition) that injects cross-domain requirement noise into the core scenario and compares monolithic and CAAF responses.
5.1 Experimental Setup
We extended the L3 AD scenario with two noise categories:
  1. 1.
    Information Volume Noise: Irrelevant constraints on CPU core scheduling and Vehicle-to-Everything (V2X) 5G video compression bandwidth.
  2. 2.
    Business Conflict Noise: A “VIP Sleep Mode” constraint mandating that longitudinal braking jerk must not exceed $1.5\,\text{m/s}^{3}$ to avoid waking a rear-seat passenger.
The monolithic baseline used GPT-4o at temperature==0.0 (n=20n{=}20 per condition). CAAF used the all-GPT-4o-mini configuration (executor temp==0.7, reviewer temp==0.0, n=20n{=}20 per condition).
5.2 Monolithic Failure: Systematic Safety-Priority Collapse
Across all 20 trials of the Safety-Only (clean) Product Requirements Document (PRD), the monolithic model converged to a mean speed of $57.9$ km/h, with 18/20 runs violating rear safety and 2/20 violating forward safety. No run passed UAI verification (0/20). The dominant failure mode is systematic prioritization of the forward safety constraint, silently violating the rear safety invariant without flagging the paradox.
Adding VIP Sleep Mode, thermal protection, and compute allocation noise (the noise condition) shifted the failure distribution: mean speed rose to $63.7$ km/h, with 14/20 runs violating rear safety and 6/20 violating forward safety, but zero outputs passed UAI verification (0/20). Across all 40 trials in both conditions, no monolithic output achieved physical compliance.
Table 4: Context Rot Benchmark Results (n=20n{=}20 per condition)
System Test Condition Paradox Detected Failure Distribution Mono GPT-4o (t=0) Clean (safety-only) 0/20 Rear-Viol: 18, Fwd-Viol: 2 Mono GPT-4o (t=0) Noise (VIP+thermal) 0/20 Rear-Viol: 14, Fwd-Viol: 6 CAAF-all-mini Clean (safety-only) 20/20 CAAF-all-mini Noise (VIP+thermal) 20/20
The noise condition does not improve the monolithic failure rate (already 0/20 in the clean condition) but shifts the failure distribution—forward safety violations increase from 2/20 to 6/20—revealing that attention is stochastically redistributed across competing requirements under noise.
5.3 CAAF Defense: Context Firewall via RAD
When processed through CAAF, the Orchestrator’s RAD established a Context Firewall. The Kinematics_Executor received only kinematics-relevant context; the VIP Sleep Mode and jerk constraint were routed to a separate comfort analysis node. In all 20 runs of both conditions (40 total), the UAI immediately detected the physical paradox—no speed value simultaneously satisfies both hard constraints—and halted with FAILED_PARADOX. CAAF’s detection is invariant to requirement noise because the UAI evaluates physical assertions independently of whatever business-context tokens the executor received.
6 Empirical Variant II: Stochastic Oscillation vs. Monotonic Convergence
Having demonstrated CAAF’s interception capability, we evaluated its convergence behavior under a naive reflection baseline (n=20n{=}20 per condition).
6.1 The Seesaw Effect of Naive Reflection
We placed a monolithic LLM (GPT-4o-mini, temperature==0.7) in an AutoGPT-style (Wu et al., 2023) naive reflection loop: upon each UAI failure, the error message was appended to the prompt as feedback, with no State Locking applied. Across n=20n{=}20 independent trials (5 iterations each), the model experienced severe Stochastic Oscillation in all 20/20 runs—zero convergence.
Table 5: Stochastic Oscillation vs. Monotonic Convergence — Aggregate (n=20n{=}20)
Iteration Dominant Speed (km/h) Freq. Other Speeds Observed CAAF — Status 1 55 (fails Rear) 14/20 84 (4), 0 (2) Paradox detected (halted) 2 84 (fails Fwd) 15/20 55 (5) 3 55 (fails Rear) 15/20 84 (5) 4 84 (fails Fwd) 13/20 55 (5), 70 (1), 110 (1) 5 55 (fails Rear) 14/20 84 (5), 90 (1)
Aggregate over n=20n{=}20 naive reflection runs (5 iterations each). 0/20 converged. The dominant pattern is $5\leftrightarrow$84 oscillation, but 20% of runs include outlier speeds (0, 70, 90, 110 km/h)—attempts at compromise values that still violate at least one constraint. CAAF (n=20n{=}20): paradox detected on the first pass, halting in all 20/20 runs.
The oscillation pattern is structurally determined: when the model attempts to fix the forward safety constraint (lowering speed to $\approx$55 km/h), it immediately violates the rear safety deceleration limit. When corrected to raise speed to $\approx$84 km/h, it violates forward safety. Occasional outlier speeds (0, 70, 90, 110 km/h) represent the model’s stochastic attempts at compromise, but all remain physically invalid. Without a mechanism to lock confirmed constraint boundaries and escalate the authorization decision, the model oscillates indefinitely.
6.2 CAAF: Deterministic Termination vs. Iterative Oscillation
CAAF exits on the first pass in all 20/20 runs (FAILED_PARADOX), requiring zero iterative guessing. The contrast is structural: on paradox inputs, CAAF halts in the first global review without iterative magnitude search—the Reviewer evaluates the constraint set, determines from deterministic UAI failures that no valid value exists, and escalates immediately. This deterministic termination property is what distinguishes CAAF from reflection-based approaches: rather than iterating toward a solution that cannot exist, CAAF declares infeasibility via the Reviewer’s root-cause analysis over deterministic UAI failures and invokes the Strategic Negotiation process (Section 3.5) to resolve the deadlock through authorized constraint relaxation.
7 Cross-Domain Validation: Pharmaceutical Flow Reactor Paradox
To validate CAAF’s domain-agnosticism claim beyond automotive engineering, we constructed a second paradox benchmark in the pharmaceutical process engineering domain. This benchmark is structurally more demanding than the AD case: it involves 7 simultaneous constraints with nonlinear (Arrhenius-exponential) interactions, a 3-way minimal unsatisfiable subset, a multi-layer DAG topology, and 4 constraints that independently PASS—enabling a rigorous test of State Locking.
7.1 Constraint Setup
A pharmaceutical process engineer must design operating parameters for a continuous flow microreactor producing a drug intermediate via a first-order reaction with a competing side reaction. The reaction follows Arrhenius kinetics:
$k(T)=A\cdot\exp\!\left(\frac{-E_{a}}{R\cdot(T+273.15)}\right),\quad A=2.5\times 10^{8}\;\text{s}^{-1},\;E_{a}=72{,}000\;\text{J/mol}$ (4)
with conversion $X=1-\exp(-k\cdot\tau)$ and impurity fraction $I=\alpha\cdot k^{2}\cdot\tau$ ($\alpha=0.35$ s). Seven constraints must be simultaneously satisfied:
Table 6: Pharmaceutical Flow Reactor Constraints. ICH = International Council for Harmonisation (pharmaceutical regulatory guidelines).
ID Constraint Formula Source C1 Conversion $X\geq 0.95$ ICH Q6A regulatory C2 Impurity $I\leq 0.02$ ICH Q3A guideline C3 Temperature $T\leq 150\,^{\circ}$C Thermal decomposition C4 Residence time $\tau\leq 120$ s Process stability C5 Production rate $\geq 5.0$ kg/day Scale-up requirement C6 Thermal safety $Q_{\text{gen}}\leq Q_{\text{cool}}$ Runaway prevention C7 Pressure drop $\Delta P\leq 15$ bar Equipment rating
The Irreconcilable Paradox. C1 and C2 interact through the shared nonlinear variable $k(T)$:
  • C1 requires $k\cdot\tau\geq-\ln(0.05)\approx 3.0$, pushing $\tau$ upward.
  • C2 requires $\alpha\cdot k^{2}\cdot\tau\leq 0.02$, capping $k^{2}\cdot\tau$.
  • Combining: $k\leq\frac{I_{\max}/\alpha}{-\ln(1-X_{\min})}\approx 0.01908\;\text{s}^{-1}$, forcing $\tau\geq 157.1$ s.
  • C4 caps $\tau\leq 120$ s. The gap is 37.1 s—no valid $\tau$ exists.
[…]
7.2 Experimental Results
We evaluated four conditions (n=20n{=}20 each) using GPT-4o-mini, adding a Mono+UAI (ReAct (Yao et al., 2023)) ablation condition to isolate the contribution of the deterministic UAI tool:
Table 7: Pharmaceutical Flow Reactor Paradox Results (n=20n{=}20 per condition)
# System Hint $n$ Correct% Failure Mode Distribution 1 Monolithic GPT-4o-mini × 20 0% Impurity-Viol: 10, Conv-Viol: 7, JSON: 3 2 Monolithic GPT-4o-mini 20 65% Impurity-Viol: 6, JSON: 1 3 CAAF-all-mini × 20 100% Paradox detected: 20/20 4 Mono + UAI (ReAct) × 20 95% Impurity-Viol: 1
Figure 10: Pharmaceutical flow reactor paradox zone in the $(T,\tau)$ plane. The C1 conversion boundary (green) requires $\tau$ above the curve; the C2 impurity boundary (red) requires $\tau$ below it. Their intersection forms the C1+C2 feasible region (green wedge), whose minimum-$\tau$ point—the feasible boundary—is at $T=98.6\,^{\circ}$C, $\tau=157$ s. The C4 residence-time limit (blue dashed) caps $\tau$ at 120 s (blue band below). The two feasible regions do not overlap: the wedge apex sits 37 s above the C4 ceiling (red arrow), constituting an empty solution space.
Figure 11: Failure mode distribution across six experimental conditions (n=20n{=}20 each) in the pharmaceutical domain. Baselines exhibit diverse failure modes dominated by impurity and conversion violations, with the sequential baseline additionally exhibiting a high JSON-failure rate (6/20) after multiple retry cycles. CAAF achieves perfect paradox detection; Mono+UAI’s single failure (1/20) is an impurity constraint omission that RAD’s atomic decomposition is designed to prevent (not directly ablated; see Finding 8).
7.3 Analysis
Finding 6 — Structural Complexity Does Not Degrade CAAF. Despite the 7-constraint, nonlinear, 3-way paradox structure, CAAF achieves 100% paradox detection (20/20), replicating the AD domain result. The DAG decomposition correctly routes temperature and residence time computations through dependent constraint nodes, and State Locking confirms that C3, C5, C6, and C7 lock to PASS throughout iteration.
Finding 7 — Hint Fragility Scales with Complexity. In the AD domain (2-constraint paradox), the hint condition achieved 100% detection. In the pharma domain (3-way paradox with nonlinear interactions), hint-aided detection drops to 65%. This confirms that prompt engineering reliability degrades as constraint complexity increases—precisely the regime where architectural enforcement becomes most critical.
Finding 8 — Mono+UAI Ablation: UAI as the Core Contribution. The Mono+UAI condition strips away RAD and Structured Semantic Gradients, giving a single monolithic LLM direct access to the UAI tool via ReAct-style function calling. At 95% (19/20), this ablation demonstrates that the deterministic UAI assertion engine provides the majority of CAAF’s value. The single failure (1/20) was an impurity constraint violation: the model selected $T{=}130\,^{\circ}$C, $\tau{=}60$ s but failed to detect that the resulting impurity fraction exceeded the 2% limit—an omission that RAD’s atomic decomposition is designed to prevent by assigning impurity verification to a dedicated node, though we did not ablate this specific combination (Mono+UAI+RAD).
This result is consistent with CAAF’s design philosophy: the three architectural pillars address complementary failure surfaces:
  • UAI (Harness as an Asset) provides the deterministic constraint boundary—the core reliability mechanism.
  • RAD ensures input fidelity at scale by decomposing complex specifications into atomic, context-firewalled subtasks, preventing context attention decay.
  • Structured Semantic Gradients with State Locking guarantee monotonic convergence across iterations, preventing oscillation between constraint fixes.
In the current 7-constraint benchmark, a capable model with UAI access can reason through the constraint interactions in a single pass. We hypothesize that RAD’s contribution will become critical at higher constraint counts (15+) and longer specifications, where context window limitations degrade single-pass reasoning. This represents a key axis for future experimental validation.
8 Alternative Architecture Baselines
To demonstrate that CAAF’s reliability derives from its specific architectural properties (UAI + State Locking) rather than from multi-agent orchestration per se, we evaluated two alternative multi-agent architectures that lack deterministic constraint enforcement.
8.1 Multi-Agent Debate Baseline
Following Du et al. (2024), we implemented a debate architecture: two LLM agents independently generate solutions, then engage in 3 rounds of cross-critique. No UAI or State Locking is applied; agents rely on mutual persuasion to converge. We evaluated on both the AD and pharmaceutical domains (n=20n{=}20 each).
8.2 Sequential Checker Baseline
We implemented a sequential pipeline: a primary LLM executor generates a solution, then a second LLM instance acts as a checker, critiquing the output. Up to 3 retry cycles are permitted. The checker uses natural language judgment (no UAI assertions).
8.3 Results
Table 8: Alternative Architecture Baselines (n=20n{=}20 per condition)
System Domain $n$ Correct% Failure Mode Distribution Debate (2 agents, 3 rounds) AD 20 0% Fwd-Viol: 7, Rear-Viol: 13 Debate (2 agents, 3 rounds) Pharma 20 0% Impurity-Viol: 15, Conv-Viol: 4, JSON: 1 Sequential (exec + checker, 3 retries) AD 20 0% Forced: 9, Fwd-Viol: 7, Rear-Viol: 4 Sequential (exec + checker, 3 retries) Pharma 20 0% Impurity-Viol: 7, Conv-Viol: 5, Dual: 2, JSON: 6 CAAF-all-mini (no hint) AD 30 100% Paradox detected: 30/30 CAAF-all-mini (no hint) Pharma 20 100% Paradox detected: 20/20 Mono + UAI (ReAct) AD 20 85% Fwd-Viol: 3 Mono + UAI (ReAct) Pharma 20 95% Impurity-Viol: 1
Finding 9 — Multi-Agent Orchestration Alone is Insufficient. Both the debate and sequential baselines achieve 0% paradox detection across all 80 trials (40 AD + 40 pharma). The debate architecture exhibits a notable failure mode: agents converge on a socially negotiated “consensus” that silently violates one constraint, producing False Consensus—in the pharma domain, 15/20 debate trials converge on impurity-violating solutions where neither agent detects the three-way paradox. The sequential baseline shows even higher noise: 6/20 pharma trials produce unparseable JSON after 3 retries, and the LLM checker approves constraint-violating solutions despite all containing physical violations. Meanwhile, CAAF achieves 100% across both domains. The Mono+UAI ablation achieves 85% on AD and 95% on pharma—substantially better than all baselines, but not perfect, confirming that the UAI assertion engine is the dominant contributor while RAD and State Locking close the remaining gap.
These results confirm that CAAF’s reliability derives from its UAI assertion engine and State Locking mechanism, not from the general principle of multi-agent decomposition. Without a deterministic external validator, adding more agents or more rounds of critique did not bridge the controllability gap at the depths tested (3 debate rounds, 3 checker retries).
Figure 12: Paradox detection rate across architectures and domains. All baselines—monolithic, debate, and sequential—achieve 0% detection in both domains. Only CAAF achieves 100% across both. The Mono+UAI ablation (85% AD, 95% pharma) demonstrates that UAI provides the majority of value, while RAD and State Locking close the remaining gap to 100%.
9 Economics and Organizational Dynamics
9.1 Paradigm Shift: From One-Shot Heuristics to Expansion-to-Convergence
The adoption of CAAF necessitates a transition in user expectations from the “One-Shot” fallacy to an Expansion-to-Convergence (E2C) workflow.
In the prevailing “One-Shot” approach, users treat LLMs as high-capability search engines, expecting a flawless engineering specification from a single inference. While initial outputs often appear syntactically impressive, they frequently harbor latent physical contradictions that surface only during downstream testing—at far greater cost than the original generation. CAAF accepts initial output variance and treats the first generation as a “coarse expansion” of possibilities, then systematically prunes this draft via Structured Semantic Gradients until it aligns with physical reality. The cost of convergence is paid in cheap inference tokens, not in expensive engineering labor.
9.2 The Economics of Certainty: Redefining Total Cost of Ownership
We formalize the Total Cost of Ownership (TCO) for AI-generated engineering artifacts:
$\text{TCO}=(N_{\text{loops}}\times C_{\text{inf}})+(P_{\text{fail}}\times C_{\text{cat}})+C_{\text{HITL}}$ (5)
where $N_{\text{loops}}$ = internal inference cycles, $C_{\text{inf}}$ = per-cycle inference cost, $P_{\text{fail}}$ = probability of an undetected defect reaching production, $C_{\text{cat}}$ = catastrophic failure cost (recall or redesign), and $C_{\text{HITL}}$ = human debugging labor cost.
Monolithic deployments minimize $N_{\text{loops}}$ (fast single-shot generation) but accept a non-zero $P_{\text{fail}}$ and incur substantial $C_{\text{HITL}}$ as engineers diagnose “compliant hallucinations.” CAAF executes a Compute-for-Risk Arbitrage: deliberately inflating $N_{\text{loops}}$ through relentless UAI validation drives $P_{\text{fail}}$ toward zero on all modeled constraints—empirically observed at 100% detection across both reported benchmarks—while reducing $C_{\text{HITL}}$ from “hunting for hidden paradoxes” to “executive decision authorization.”
Empirical TCO Analysis (L3 AD Paradox Benchmark): API costs were directly measured across n=30n{=}30 trials per condition. Enterprise pricing (illustrative reference rates): GPT-4o at \$2.50/\$10.00 per 1M tokens (input/output); GPT-4o-mini at \$0.15/\$0.60 per 1M tokens.
Table 9: Empirical TCO Comparison
Metric Mono GPT-4o Mono GPT-4o-mini CAAF-all-mini API cost per run (measured) \$0.0145 \$0.0006 \$0.0027 Paradox detection rate 0% 0% 100% Artifact physically verified? No No Yes Estimated HITL cost (per failure) \$50.00 \$50.00 \$5.00 Effective cost per correct artifact $\infty$ $\infty$ \$0.0027
[…]
HITL cost basis: \$50/incident == $\sim$20 min of senior systems engineer time to diagnose a “compliant hallucination”; \$5/incident == $\sim$2 min of executive review of a pre-computed boundary report. These are conservative engineering-labor estimates for illustration; actual values are domain- and organization-specific.
Scaling Behavior: For tasks with $S$ independent constraints, monolithic joint success probability decays as $P_{\text{success}}=p^{S}$ (e.g., p=0.85p{=}0.85 per constraint). CAAF’s constraint-wise decomposition breaks this joint-failure dependency, providing approximately linear cost scaling. Table 10 projects the theoretical TCO differential as constraint complexity increases:
Table 10: Theoretical TCO Scaling (Illustrative Projection)
[…]
These projections apply the empirical per-run API costs and the HITL cost estimates above across varying constraint counts. They are theoretical projections intended to illustrate scaling behavior, not empirical measurements at each complexity level. External UAI simulation latency (e.g., FEA solvers) would shift the bottleneck from human cognition to machine compute, reinforcing the economic thesis.
Figure 13: Cost vs. reliability for the no-hint condition (n=30n{=}30). CAAF-all-mini (green) achieves 100% paradox detection at \$0.0027/run, dominating the efficiency frontier. Both monolithic baselines (GPT-4o at \$0.0145/run and GPT-4o-mini at \$0.0006/run) achieve 0% detection, yielding infinite effective cost per correct artifact.
9.3 The AI as Collaborative Decision Arbiter
Post-mortem analyses of catastrophic engineering failures—including the space shuttle Challenger O-ring incident and the Boeing 737 MAX Maneuvering Characteristics Augmentation System (MCAS) design defects—have identified systemic organizational factors as significant contributors alongside technical causes (Vaughan, 1996). In these cases, engineers possessed relevant technical knowledge but were unable to prevent the erosion of safety margins under schedule and managerial pressure. The common pattern is not ignorance of risk but institutional failure to preserve engineering red lines when they conflict with delivery timelines.
CAAF structurally addresses this behavioral vulnerability. Consider a scenario where a project manager requests that a software team “bypass the safety warning” to meet a deployment deadline. A monolithic LLM may comply. CAAF does not, by construction: if the artifact violates the frozen Harness, the UAI emits [HARD FAILURE]. To proceed, the team must engage the Strategic Negotiation process (Section 3.5), creating a formal, timestamped decision record. CAAF is designed to convert transient project pressures into durable institutional knowledge, conditional on the integrity of the Harness Registry itself (see Section 11.2).
9.4 A Constraint Satisfaction Layer for LLM-Generated Artifacts
The empirical results in this paper span two domains—L3 AD requirements engineering (Section 4) and pharmaceutical continuous flow reactor design (Section 7). We argue that the problem CAAF addresses is not limited to these, and offer the framing below as a hypothesis motivated by the architecture—one whose empirical scope remains bounded by the two benchmarks reported here. We conjecture that workflows of this shape—an LLM generating a complex artifact subject to formally verifiable constraints—share the same structural risk surface demonstrated in Sections 4–8: sycophantic compliance, context rot, stochastic oscillation, and silent constraint override; broader empirical validation of this conjecture remains future work. The pattern is general: human intent expressed in natural language $\rightarrow$ LLM-generated artifact $\rightarrow$ formal correctness criteria that must all be satisfied simultaneously $\rightarrow$ potential contradictions among criteria that must be detected, not masked.
CAAF’s domain-specific component is exclusively the Harness Registry and its UAI validators. The RAD decomposition engine, convergence control loop, State Locking mechanism, and Strategic Negotiation protocol are designed to be domain-agnostic. Replacing the YAML Harness with domain-appropriate assertions is the intended extension mechanism, and we conjecture that the convergence behavior observed on the two benchmarks carries over to any domain where constraints can be formalized as executable predicates and composed into a DAG with bounded branching factor. The UAI evaluation mechanism is identical whether the assertion computes a braking distance, a thermodynamic equilibrium, a regulatory compliance check, or a financial risk exposure—but empirical validation beyond physical-law constraints remains future work.
We identify four categories of formally verifiable constraints, each mapping to distinct industries:
  • Physical Laws (kinematics, thermodynamics, material science): Bio-pharma process design—temperature ramp stability can be formalized as thermodynamic assertions; UAI would evaluate whether proposed cooling schedules violate phase transition boundaries. Structural engineering—load-bearing calculations can be expressed as Finite Element Analysis (FEA)-derived assertions; existing simulation tools can be integrated via MCP as UAI validators.
  • Regulatory and Legal Rules (ISO standards, FDA regulations, GDPR, building codes): Clinical trial protocol design—sample size, exclusion criteria, and informed consent requirements are formalizable constraints that frequently conflict with practical constraints (budget, timeline, patient availability). CAAF’s paradox detection and Strategic Negotiation are directly applicable: when ethical constraints conflict with statistical power requirements, the system surfaces the trade-off with quantified impact rather than silently relaxing either.
  • Mathematical and Logical Rules (type systems, formal specifications, financial models): Financial risk modeling—regulatory capital requirements (e.g., Basel III/IV) impose simultaneous constraints on leverage ratios, liquidity coverage, and risk-weighted assets. These constraints are mathematically expressible and frequently tension against business objectives.
  • Infrastructure and Operational SLAs (latency, availability, cost budgets): Cloud Infrastructure as Code—Kubernetes deployment configurations must simultaneously satisfy security policies (network isolation, pod security standards), SLAs (latency << 100ms, replicas $\geq$ 3), and cost budgets. These three constraint categories are individually formalizable and routinely conflict: high availability requires more replicas, which violates cost constraints. Current tools (OPA, Terraform plan) perform post-hoc policy checks but do not provide convergence or paradox detection.
Embodied AI and Vision-Language-Action (VLA) Models. A particularly compelling application frontier is the emerging class of Vision-Language-Action models (e.g., RT-2 (Brohan et al., 2023), Octo (Octo Model Team et al., 2024)) for robotics and autonomous vehicles. VLA models generate action plans from visual perception and natural language instructions, but remain fundamentally probabilistic—they provide no formal guarantee that a proposed plan satisfies safety constraints. Current safety mechanisms in embodied AI operate primarily at the control layer (millisecond-level): Control Barrier Functions (CBFs) enforce real-time safety boundaries, action-space clipping limits individual joint ranges, and emergency-stop monitors provide a last resort. However, at the planning layer (second-level)—where VLA models decide what to do—there is no equivalent of a constraint satisfaction guarantee: if a VLA proposes a manipulation plan that simultaneously violates gripper force limits, collision boundaries, and task-completion deadlines, no existing system detects the multi-constraint conflict, converges toward a feasible plan, or identifies when the constraint set is irreconcilable.
CAAF fills this gap as a planning-layer constraint satisfaction layer that complements control-layer safety mechanisms. Each VLA-proposed action plan can be validated against a Harness encoding physical safety constraints (joint torque limits, collision boundaries), task constraints (completion time, precision tolerances), and human-safety constraints (approach speed, handover force). When constraints conflict—for instance, when a manipulation task requires a trajectory that is fast enough to meet a deadline but too fast for safe human proximity—CAAF’s paradox detection surfaces the trade-off with quantified options rather than silently selecting a compromised plan. This positions CAAF and CBFs as complementary layers: CAAF ensures the plan is correct before execution; CBFs ensure execution stays safe during execution.
In each case across the domains above, the distinguishing feature of CAAF is not merely constraint checking (which existing tools already provide) but the combination of convergence toward satisfaction, paradox detection when the constraint set is irreconcilable, and structured negotiation with quantified trade-offs. No existing framework provides all three capabilities in a unified architecture.
We provide empirical validation for the first category (Physical Laws) in Section 7. Validation across the remaining categories constitutes a primary direction for future research (Section 11.6).
Enterprise Offline Deployment. The experimental results (Finding 5, Section 4.2, Table 2) empirically establish that CAAF’s reliability does not depend on frontier closed-source model capability. Two open-weight families at two distinct parameter scales—Cohere Command-R7B (7B, structured-output fine-tuned) and Google Gemma-3-12B-IT (12B, general-purpose)—each achieve 100% on both the AD PASS-path benchmark (20/20 SUCCESS) and the pharma 3-way paradox benchmark (20/20 paradox detection), exactly matching the GPT-4o-mini reference configuration. For regulated industries where data-residency constraints prohibit transmission to external APIs (automotive type approval, ISO 13485 medical devices, IEC 62443 industrial control systems (International Organization for Standardization, 2018)), a fully offline CAAF instance using one of these open-weight models can in principle be certified as part of the engineering toolchain without sacrificing the convergence guarantee. The Harness Registry and UAI remain deterministic regardless of model host, and the convergence behavior is architectural rather than API-dependent. This positions CAAF as a practical path to AI-augmented requirements engineering in environments currently closed to cloud-dependent LLM tooling, with the one engineering prerequisite surfaced by our replication study: the selected open-weight model must either (a) be fine-tuned for structured / tool-use output, or (b) be large enough ($\geq$12B) that general instruction-following is sufficient to emit schema-faithful JSON for the Orchestrator’s decomposition step.
10 Related Work
10.1 LLM Agent Frameworks
Prompt Engineering treats the LLM as a black-box stochastic optimizer, refining inputs to maximize the probability of a desirable output distribution. CAAF explicitly supersedes this approach: rather than optimizing the input to a single inference, it applies a formal cybernetic control loop to the output across multiple iterations. The Structured Semantic Gradient is fundamentally different from a prompt improvement—it is a directed corrective signal anchored in physical-space measurements.
[…]
LangGraph (LangChain, Inc., 2024) excels at durable, stateful graph-based orchestration. Its architectural limitation is Node Logic Leakage: without manually implemented guards on each node, unconstrained nodes can corrupt downstream context. CAAF automates this defense via the Constraint Awareness Layer and UAI, eliminating the need for developers to anticipate every possible cross-domain constraint interaction.
AutoGen (Wu et al., 2023) and OpenAI Swarm enable multi-agent actor handoffs and social debate. In complex physical environments, debate without a physical grounding interface is susceptible to Stochastic Oscillation (Section 6) and False Consensus (Section 8). CAAF replaces qualitative social handoffs with unarguable semantic gradients mechanically mapped to UAI results.
Multiagent Debate (Du et al., 2024) has demonstrated improvements in factuality and reasoning for open-domain tasks. However, in constraint satisfaction scenarios with irreconcilable physical paradoxes, agents may collectively converge on a hallucinated solution (e.g., “assume the rain has stopped”) rather than declaring a deadlock. UAI Hard Assertions prevent this by providing an incorruptible external verdict.
AgentForge (Anbar Jafari et al., 2026a) represents a recent class of lightweight modular agent frameworks that compose skills as a DAG with YAML-based agent configuration, sharing CAAF’s preference for declarative, composable orchestration over monolithic prompt engineering. Their focus is developer ergonomics and reduced integration overhead; CAAF complements this direction by adding a deterministic external validator (UAI) and a convergence loop on top of the same compositional primitives, addressing the safety-critical regime where “correct most of the time” is insufficient.
10.2 LLM Guardrails and Safety
Constitutional AI (Bai et al., 2022) encodes safety principles into model behavior through training-time feedback. This internalizes constraints into model weights. CAAF externalizes constraints into UAI assertions: a model weight can drift or be overridden by fine-tuning; a Python assertion cannot. For regulatory environments requiring immutable, auditable constraint records, CAAF’s external approach provides stronger guarantees.
NeMo Guardrails (Rebedea et al., 2023) provides programmable rails for LLM applications, primarily targeting semantic safety (topical constraints, factual accuracy). CAAF targets physical safety: the UAI evaluates mathematical assertions over numerical engineering parameters. NeMo provides Pass/Fail intercept; CAAF provides intercept plus Convergence.
Guardrails.ai (Guardrails AI, 2023) offers structured output validation and automatic re-asking. Like NeMo, it focuses on single-shot output structure rather than multi-step constraint convergence across a topological dependency graph.
Consequence-Aware Agentic AI (CA2I) (Anbar Jafari et al., 2026b) is concurrent work that introduces a pre-execution outcome-assessment module for responsible LLM agents, sharing CAAF’s premise that constraint violations should be caught before the agent acts rather than detected post-hoc. The two designs are complementary along orthogonal axes: CA2I evaluates the projected consequences of a proposed action against high-level responsibility criteria, whereas CAAF formalizes domain invariants as an executable Harness Registry and uses a deterministic UAI plus monotonic-convergence loop to drive the agent toward a verifiable artifact. Combining the two—using consequence assessment as an additional UAI clause inside the Harness—is a natural integration path.
10.3 Textual Gradients and Optimization
TextGrad (Yuksekgonul et al., 2024) applies the backpropagation metaphor to text, using an LLM evaluator to provide textual gradients for offline prompt and code optimization. CAAF applies a related concept to real-time, zero-shot task execution: the UAI provides a deterministic error trace (analogous to a loss signal), and State Locking ensures that gradient application is monotonically improving. TextGrad’s gradients are fully LLM-generated; CAAF’s gradients are physically anchored by the UAI, making them partially deterministic.
10.4 Neuro-Symbolic and Formal Methods
CAAF instantiates a practical Neuro-Symbolic architecture (Schick et al., 2023): the LLM component handles unstructured natural language, heuristic search, and strategic communication, while symbolic components (Python assertion engine, simulation tools, TLA+ verifiers) provide rigorous mathematical evaluation. This division of labor is analogous to the System 1/System 2 distinction in cognitive science—fast, associative reasoning augmented by slow, deliberate verification.
SMT Solvers (Z3, CVC4). Satisfiability-modulo-theories solvers (de Moura and Bjørner, 2008; Barrett et al., 2011) are the gold standard for deciding constraint satisfaction over formal theories (linear arithmetic, bit-vectors, arrays) and could, in principle, subsume CAAF’s paradox-detection role. The key differentiator is the input boundary: Z3 requires the full problem to already be expressed in SMT-LIB (or an equivalent formal language), whereas CAAF accepts natural-language requirements authored by domain engineers and carries the burden of lifting those requirements into executable assertions through the Orchestrator/RAD pipeline. The two approaches are complementary rather than competing: within an individual UAI validator, an SMT solver is a legitimate implementation choice (e.g., a torque_budget rule could dispatch to Z3 for a mixed-integer linear subproblem), and we view such embeddings as natural extensions. What CAAF adds on top of any such solver is (i) the natural-language-to-assertion lift, (ii) monotonic convergence with State Locking rather than a single yes/no verdict, and (iii) a paradox-surfacing protocol that returns quantified trade-offs to a human operator, rather than an unsat core that requires a formal-methods specialist to interpret.
Constrained / Guided Decoding. Systems such as Guidance (Microsoft Research, 2023) and Outlines (Willard and Louf, 2023) enforce constraints at the token level by masking the decoder’s output distribution so that generated text is guaranteed to conform to a regular expression, context-free grammar, or JSON schema. This is extremely effective for structural constraints (valid JSON, valid SQL, well-typed function calls) but does not address semantic-physical constraints: a perfectly schema-compliant JSON document can still specify $v_{t5}{=}120\,\text{km/h}$ under heavy rain, violate an Arrhenius impurity bound, or trip a forward-safety rule. CAAF operates at a strictly higher layer—the artifact layer, after the full candidate has been generated and parsed—and verifies constraints that require arbitrary computation (Arrhenius kinetics, kinematic equations, regulatory lookups) rather than only finite-state language membership. The two are compositional: constrained decoding can be used to guarantee that each Executor emits a parseable Harness-conformant artifact in the first place, after which CAAF’s UAI performs the semantic validation and gradient-based revision that constrained decoding cannot express.
Harness Engineering (Böckeler, 2024) established patterns for managing LLM interaction environments to improve reliability in developer productivity tools. CAAF formalizes these patterns into a rigorous framework with guaranteed convergence properties, shifting the application domain from coding assistance (Agent-as-Tool) to safety-critical engineering governance (Agent-as-Controller).
10.5 Comparative Summary
Table 11 contrasts CAAF with prior agent and constraint-checking frameworks along the four capabilities highlighted throughout this paper: deterministic constraint checking, convergence behavior, paradox detection on irreconcilable inputs, and structured negotiation of trade-offs. To our knowledge, no prior system provides all four in a unified architecture.
Table 11: Comparative summary of frameworks. Dashes (—) indicate absence. CAAF is the only framework providing all four capabilities in a unified architecture.
Framework Constraint Check Convergence Paradox Detection Structured Negotiation AutoGPT / BabyAGI LangGraph (LangChain, Inc., 2024) Manual guards Manual trapping AutoGen (Wu et al., 2023) Qual. reflection NeMo / Guardrails.ai Pass/fail OPA (CNCF, 2019) Deterministic TextGrad (Yuksekgonul et al., 2024) LLM-judged LLM gradient Z3 / SMT (de Moura and Bjørner, 2008) Formal (SMT-LIB in) Decision procedure unsat core Outlines / Guidance Token-level (schema) CAAF (ours) Deterministic (UAI) State Lock + Boundary Grad. Topological RCA Quantified Trade-offs + HITL
11 Limitations and Future Work
11.1 Orchestrator Context Horizon
CAAF’s RAD mitigates structural hallucination via deterministic parsing of Harness metadata. However, in hyper-complex industrial systems (e.g., Very-Large-Scale Integration (VLSI) chip design, macro-scale supply chains), even the Orchestrator may eventually face context saturation when managing thousands of interdependent sub-graphs.
We propose Recursive Sub-Graph Partitioning (RSP) as the mitigation strategy. The master Orchestrator applies a topological clustering algorithm (e.g., Louvain or Spectral Clustering) to divide the global graph into semantically cohesive sub-domains, each managed by a dedicated Sub-Orchestrator. Global convergence is achieved via a recursive merge-and-audit step at domain interfaces. This hierarchical approach scales CAAF logic to systems of arbitrary complexity.
11.2 Harness Fidelity Bottleneck
CAAF’s determinism guarantees are bounded by Harness completeness. For domains where physical invariants are poorly understood or mathematically unformalized (e.g., emergent market sentiment, aesthetic design preferences), the UAI cannot provide grounding signals. In such domains, CAAF’s structural advantages are reduced, and the system degrades gracefully toward a structured multi-agent orchestration framework without convergence guarantees.
11.3 Convergence Guarantees
[…]
11.4 Computational Overhead and Latency
[…]
11.5 Two-Domain Validation Scope
The empirical claims of this paper rest on two benchmark domains: L3 autonomous driving requirements engineering (Section 4) and pharmaceutical continuous flow reactor design (Section 7). The two benchmarks are structurally complementary: the AD paradox involves 2 scalar constraints with a linear gap, while the pharma paradox involves 7 constraints with nonlinear (Arrhenius-exponential) interactions and a 3-way minimal unsatisfiable subset. This structural diversity strengthens the domain-agnosticism claim. Nevertheless, both benchmarks have analytically transparent ground truth. Future work will validate CAAF on domains where ground truth is less analytically transparent (e.g., financial risk modeling, supply chain optimization) and on constraint systems with even higher dimensionality (15+ constraints).
11.6 Broader Multi-Domain Empirical Validation
Extend the empirical evaluation to additional constraint-paradox benchmarks beyond the two domains validated here (e.g., financial regulatory compliance, supply chain logistics, energy grid optimization). Each benchmark requires a formalized Harness with verifiable ground truth. Validating CAAF on constraint systems where single-pass reasoning degrades (15+ constraints, multi-page specifications) would substantiate the hypothesis that RAD’s contribution becomes critical at scale—a prediction arising from the Mono+UAI ablation results (Section 7).
11.7 Automated Multi-Objective Constraint Negotiation
Currently, constraint relaxation requires human-in-the-loop authorization. Future iterations will formalize the negotiation phase by extending the Harness schema with hierarchical variable definitions (fixed vs. negotiable) and cost attributes encoding multi-dimensional penalty functions (time, cost, safety margin, brand reputation). When the UAI detects an unresolvable paradox, a symbolic solver (e.g., SymPy-based algebraic engine) will automatically formulate the resolution as a Multi-Objective Constrained Optimization problem, presenting human operators with mathematically verified, cost-optimal relaxation paths.
12 Conclusion
As AI transitions from creative exploration to industrial governance, the primary engineering challenge shifts from maximizing raw intelligence to enforcing deterministic control. CAAF addresses this challenge by supplanting fragile probabilistic self-correction with a rigorous cybernetic feedback loop anchored in physical reality. By formalizing domain knowledge into an executable Harness and utilizing RAD to maintain context fidelity, CAAF provides a pathway toward substantially higher-reliability AI-driven engineering systems within the scope of a well-maintained Harness.
The Mono+UAI ablation (Section 7) reveals that the deterministic UAI assertion engine provides the majority of CAAF’s value, achieving 95% paradox detection even without multi-agent decomposition. The three architectural pillars address complementary failure surfaces: UAI provides deterministic constraint boundaries, RAD ensures input fidelity at scale, and Structured Semantic Gradients with State Locking guarantee monotonic convergence. We hypothesize that RAD’s contribution becomes critical at higher constraint counts where single-pass reasoning degrades—a key axis for future validation.
More broadly, we propose CAAF as a candidate architectural pattern—a constraint satisfaction layer for LLM-generated artifacts—intended to apply wherever the output of an LLM must simultaneously satisfy formally verifiable constraints. Just as type systems catch programming errors at compile time before code reaches production, CAAF catches constraint violations at generation time before engineering artifacts enter downstream processes. The domain-specific component is exclusively the Harness Registry; the convergence mechanism, paradox detection, and structured negotiation are designed to be domain-agnostic. On the two reported benchmarks the architectural behavior is identical: either the artifact satisfies all constraints, or the system formally reports which constraints are irreconcilable and presents quantified resolution options for human authorization. Extending this validation to regulatory, financial, and infrastructure-SLA constraint families is the primary direction for future work (Section 11.6). Subject to that broader validation, we view CAAF not as a domain-specific tool but as a prospective reliability layer for the emerging class of LLM-augmented engineering workflows in constraint-governed industries.
Appendix A Domain-Specific Harness Assets (AD Degradation)
The harness below is the exact production file used in the experiment (see harness/data/ad_degradation.yaml in the repository). Physics constants (e.g., vehicle_speed_kmph_t0) are injected by the UAI at assertion time and are never sent to the LLM, preventing the model from reasoning around physical boundaries.
Listing 4: AD Degradation Harness (YAML)
# OpenCAAF Domain Constraints: L3 Autonomous Driving Degradation
rules:
- id: REAR_COLLISION_PREVENTION_DECELERATION
description: "Rear Safety Constraint: To prevent rear-end collisions
during the takeover window, the velocity drop must not exceed
the deceleration limit."
target_field: "vehicle_speed_kmph_t5"
condition: "(vehicle_speed_kmph_t0 - vehicle_speed_kmph_t5) /
(transition_window_seconds * m_per_sec_to_km_per_h_factor)
<= max_deceleration_limit"
assertion: "(input.get(’vehicle_speed_kmph_t0’)
- input.get(’vehicle_speed_kmph_t5’)) /
(input.get(’transition_window_seconds’)
* input.get(’m_per_sec_to_km_per_h_factor’))
<= input.get(’max_deceleration_limit’)"
severity: "CRITICAL"
- id: FORWARD_COLLISION_PREVENTION_PERCEPTION
description: "Forward Safety Paradox: To prevent a blind forward
collision, the physical braking distance at the target speed
MUST be strictly less than the perception range."
target_field: "vehicle_speed_kmph_t5"
condition: "((vehicle_speed_kmph_t5 / m_per_sec_to_km_per_h_factor)
** 2) / (2 * road_friction_mu * g) < perception_range_limit"
assertion: "((input.get(’vehicle_speed_kmph_t5’)
/ input.get(’m_per_sec_to_km_per_h_factor’)) ** 2)
/ (2 * input.get(’road_friction_mu’) * input.get(’g’))
< input.get(’perception_range_limit’)"
severity: "FATAL"
Appendix B Multi-Agent Convergence Trace (AD Paradox)
The trace below shows the CAAF pipeline reaching FAILED_PARADOX. All roles (Orchestrator, Executor, Reviewer) use gpt-4o-mini, consistent with the all-commodity-model configuration validated in the batch experiment.
Listing 5: CAAF Convergence Trace (JSON)
"step": "Decomposition",
"role": "strategy_engine",
"model": "gpt-4o-mini",
"nodes": {
"node_1": {
"id": "node_1",
"parent_id": null,
"description": "Define safe-state speed limit from perception data",
"context_keys": [],
"expected_schema": { "perception_range_m": "float" }
"node_2": {
"id": "node_2",
"parent_id": "node_1",
"description": "Calculate stopping distance at target speed",
"context_keys": ["perception_range_m"],
"expected_schema": { "vehicle_speed_kmph_t5": "int" }
"review": {
"id": "FORWARD_COLLISION_PREVENTION_PERCEPTION",
"status": "FAIL",
"error": "Stopping distance (82m) > perception limit (30m)"
Appendix C Formal Deadlock Evidence Package
Listing 6: Formal Deadlock Evidence Package
# [SYSTEM DEADLOCK] Formal Paradox Report
Status: FAILED_PARADOX
Domain: L3 Autonomous Driving Degradation
## Conflict Summary
The system has encountered a mathematical deadlock between two
non-negotiable physical red lines:
1. REAR_COLLISION: To prevent highway rear-end collision,
target speed must be >= 84 km/h.
2. FORWARD_COLLISION: To stop within 30m perception range,
target speed must be <= 55 km/h.
## Evidence
- Attempting speed = 84 km/h: Forward Safety FAIL
(70m stopping distance > 30m vision)
- Attempting speed = 55 km/h: Rear Safety FAIL
(decel = 3.6 m/s^2 > 2.0 m/s^2 limit)
## Resolution
Human strategic authorization required.
See Strategic Resolution Menu.
Appendix D Reproducibility and Code Availability
All experiments described in this paper are fully reproducible. The complete codebase, including the OpenCAAF framework implementation, harness definitions, experiment scripts, and raw result logs, is available at:
Repository (Apache-2.0; see project README for setup instructions): https://github.com/TianbaoZhang001/OpenCAAF
How to run. Each benchmark is a self-contained Python module invoked from the project root. Setup requires an OpenAI API key in .env (OPENAI_API_KEY) and pip install -r requirements.txt. The sample size for any benchmark can be overridden via the N_TRIALS environment variable (default n=20n{=}20, except benchmark_full_experiment.py which defaults to n=30n{=}30). Example:
N_TRIALS=20 python -m OpenCAAF.demos.benchmark_ad_pass_path
Each run writes a timestamped directory under OpenCAAF/demos/logs/ containing results.json (aggregate metrics), *_runs.jsonl (per-trial records), and, for CAAF runs, a *_traces/run_NN/ subtree of per-iteration strategy plans, expert outputs, UAI feedback, and final artifacts.
Key experiment files:
  • OpenCAAF/demos/benchmark_full_experiment.py — 7-condition AD batch experiment (n=30n{=}30, Table 1)
  • OpenCAAF/demos/benchmark_ad_pass_path.py — AD PASS-path convergence on satisfiable variant (n=20n{=}20, Table 3, Figure 9)
  • OpenCAAF/demos/benchmark_pharma_reactor.py — Pharmaceutical Flow Reactor benchmark (n=20n{=}20, Table 7)
  • OpenCAAF/demos/benchmark_debate_baseline.py — Multi-agent debate baseline (n=20n{=}20, Table 8)
  • OpenCAAF/demos/benchmark_sequential_baseline.py — Sequential checker baseline (n=20n{=}20, Table 8)
  • OpenCAAF/demos/benchmark_context_rot_v2.py — Context Rot benchmark (n=20n{=}20, Table 4)
  • OpenCAAF/demos/benchmark_oscillation_v2.py — Oscillation benchmark (n=20n{=}20, Table 5)
  • OpenCAAF/demos/benchmark_pharma_pass_path.py — Pharma PASS-path variant ($\tau_{\max}{=}180$ s) — companion to the PASS-path study, retained as a harder-setting benchmark for stronger executor models (not used in the reported results)
  • OpenCAAF/harness/data/ad_degradation.yaml — AD Harness Registry (Appendix A)
  • OpenCAAF/harness/data/ad_degradation_pass.yaml — AD Harness Registry (PASS-path variant, §4.5)
  • OpenCAAF/harness/data/pharma_flow_reactor.yaml — Pharmaceutical Flow Reactor Harness Registry
  • OpenCAAF/harness/data/pharma_flow_reactor_pass.yaml — Pharmaceutical Flow Reactor Harness Registry (PASS-path variant, $\tau_{\max}{=}180$ s)
  • CAAF_paper/figures/generate_figures.py — AD figure generation scripts (including Figure 9)
  • CAAF_paper/figures/generate_pharma_figures.py — Pharma figure generation scripts
Raw result logs are archived in OpenCAAF/demos/logs/ and include per-run API traces, CAAF decision trees, and UAI assertion outputs.
The total API cost of all experiments reported in this paper—including the primary AD experiment (\$1.27), pharmaceutical reactor benchmark (\$0.14), debate and sequential baselines (\$0.34), context rot / oscillation benchmarks, and the AD PASS-path study (\$0.36)—was under \$2.20 USD total, demonstrating the reproducibility economics of the CAAF approach.